# VideoChat-R1: Enhancing Spatio-Temporal Perception via Reinforcement Fine-Tuning

### v4

@@ Abstract
Recent advancements in reinforcement learning have significantly advanced the reasoning capabilities of multimodal large language models (MLLMs). While approaches such as Group Relative Policy Optimization (GRPO) and rule-based reward mechanisms demonstrate promise in text and image domains, their application to video understanding remains limited. This paper presents a systematic exploration of Reinforcement Fine-Tuning (RFT) with GRPO for video MLLMs, aiming to enhance spatio-temporal perception while maintaining general capabilities. Our experiments reveal that RFT is highly data-efficient for task-specific improvements. Through multi-task RFT on spatio-temporal perception objectives with limited samples, we develop VideoChat-R1, a powerful video MLLM that achieves state-of-the-art performance on spatio-temporal perception tasks without sacrificing chat ability, while exhibiting emerging spatio-temporal reasoning abilities. Compared to Qwen2.5-VL-7B, VideoChat-R1 boosts performance several-fold in tasks like temporal grounding (+31.8) and object tracking (+31.2). Additionally, it significantly improves on general QA benchmarks such as VideoMME (+0.9), MVBench (+1.0), and Perception Test (+0.9). Our findings underscore the potential of RFT for specialized task enhancement of Video MLLMs. We hope our work offers valuable insights for future RL research in video MLLMs.
@@ 1 Introduction
Recent advancements in the application of reinforcement learning (RL) in the domain of large language models (LLMs) have demonstrated remarkable progress. As evidenced by OpenAI-o 1 [13], the implementation of test-time scaling strategies has shown substantial potential in enhancing LLMs’ capacity for complex reasoning. Subsequently, DeepSeek-R1-Zero [10] revealed that even without extensive supervised fine-tuning, the strategic application of a rule-based reward system for reward modeling could effectively harness reinforcement learning to unlock exceptional reasoning and cognitive capabilities in language models.
Current research endeavors have increasingly focused on replicating DeepSeek-R1’s success in multimodal large language models (MLLMs). Notably, Virgo [5] attempted to impart visual reasoning capabilities through knowledge distillation from open-source reasoning models such as DeepSeek-R1 [10], QwQ [28], and QvQ [27]. However, the predominant research direction [47, 19, 42, 3, 23, 20, 39, 44, 43, 4] emphasizes direct implementation of DeepSeek-R1’s core Group Relative Policy Optimization (GRPO) combined with its rule-based reward system to enable visual reasoning in MLLMs. This approach has primarily concentrated on enhancing performance in multimodal tasks involving mathematical reasoning with visual inputs and spatial localization challenges.
For video understanding, from the perspective of stimulating and evaluating reasoning abilities, there is no training and evaluation corpus as suitable as math problems and coding problems in the fields of text and images. Some works [46, 31, 6] conducted concurrently with ours have validated the superiority of the GRPO algorithm over supervised fine-tuning in some specific video tasks, such as temporal grounding and video question answer. However, deeper analysis and comprehensive ablation experiments focusing on video reasoning mechanisms remain underexplored. Current research gaps include systematic evaluations of the algorithm’s generalizability across diverse video-based reasoning scenarios and granular investigations into the interplay between rule-based reward systems and multimodal temporal dependencies.
@@
Compared with video reasoning, spatio-temporal perception is a direction in which it is easier to obtain training corpora and design a rule-based reward system. Taking the enhancement of the spatiotemporal perception ability of existing video MLLMs as the core, this paper systematically and comprehensively examines the effects of Reinforcement Fine-Tuning (RFT) on various video tasks, aiming to provide critical insights for future research. Our main findings are as follows.
@@
  • Through joint reinforcement fine-tuning on multiple spatio-temporal perception tasks, we construct VideoChat-R1, a powerful Video MLLM that boasts state-of-the-art spatiotemporal perception capabilities while also taking into account chat abilities. We have also discovered that training on spatio-temporal perception tasks has slightly strengthened the model’s spatio-temporal reasoning abilities. Compared with Qwen2.5-VL-7B, VideoChat-R1 achieves several times the performance improvement in spatiotemporal perception tasks such as temporal grounding (+31.8) and object track (+31.2). At the same time, it also achieves significant improvements on general QA benchmarks, such as VideoMME (+0.9), MVBench (+1.0), and Perception Test (+0.9)
  • The improvement of spatio-temporal perception ability and the preservation of the original chat capability can contribute to a more reliable and efficient video dialogue system. Our VideoChat-R1 can provide reference video segments when answering users’ questions. Meanwhile, we propose to utilize these video segments for “Clue-Perception" to further obtain more accurate answers. Our experimental results reveal the potential of the approach that enhances the model’s spatiotemporal perception ability through reinforcement learning for future research in the directions of reliable video dialogue systems and long video understanding.
@@ 2 Related work
Reinforcement Learning Enhancement for MLLMs.
Recently, works like OpenAI-o1 [13] and DeepSeek-R1 [10] have made significant breakthroughs in lifting the reasoning capabilities of large language models (LLMs) through reinforcement learning (RL). These advancements [25, 10, 26] enhance their proficiency in solving complex tasks in chains, including challenging math and coding problems. For MLLMs, many efforts [47, 19, 42, 3, 23, 20, 39, 44, 43, 4] have applied RL techniques with verifiable reward mechanisms to boost visual reasoning performance. Researches in video domain remain relatively underexplored, with only a few studies [31, 46, 6] investigating how to adapt RL-based strategies to spatiotemporal reasoning. Specifically, TimeZero [31] and R1-Omini [46] respectively demonstrate the potential of GRPO in temporal grounding and sentiment analysis. Video-R1 [6] extends GRPO to facilitate implicit temporal reasoning and achieves improvements in video spatial reasoning.
Spatio-Temporal Perception of MLLMs.
Spatio-temporal perception is one of the most core capabilities of video understanding models. Despite the significant progress that Video Multimodal Large Language Models (video MLLMs) [15, 21, 16, 32, 45, 33, 18, 12, 1] have recently made in general understanding tasks such as video question answering and captioning, MLLMs’ video performance still lag behind humans (event classical vision expert models) notably. Merlin [40] and TimeSuite [41] introduce spatio-temporal data augmentation for MLLM’s temporal abilities, at the cost of general performance. VideoChat-TPO [37] enhances fine-grained spatio-temporal perception in videos by introducing task-specific heads using substantial training costs.
@@ 3 Methodology
We first briefly review the Group Relative Policy Optimization (GRPO) [25]. Then, we demonstrate how we design and leverage the spatio-temporal rewards for GRPO to enhance video MLLMs.
3.1 Preliminary of Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [25] is a variant of Proximal Policy Optimization (PPO) [24] in reinforcement learning. By comparing groups of candidates responses directly, GRPO eliminates dependency on a critic model and significantly lowers training resources. Given an input question $q$, GRPO first generates $G$ distinct candidate responses $o=\{o_{1},\dots,o_{G}\}$ through policy sampling. A predefined reward function is used to get the corresponding scores $\{R_{1},\dots,R_{G}\}$. GRPO computes their mean and standard deviation for normalization and determines the quality of these responses:
$A_{i}=\frac{R_{i}-\mathrm{mean}(\{R_{i}\}_{i=1}^{G})}{\mathrm{std}(\{R_{i}\}_{i=1}^{G})}\text{,}$ (1)
where $A_{i}$ represents the relative quality of the $i$-th answer. GRPO encourages the model to favor better answers with a high score within the group. The final training objective also considers preventing the optimized policy $\pi_{\theta}$ from deviating far from the original MLLM parameters $\pi_{\mathrm{ref}}$ by adding a KL-divergence term $\mathrm{D}_{\mathrm{KL}}(\cdot|\cdot)$ to :
$\max_{\pi_{\theta}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}}\Big[$ $\sum_{i=1}^{G}\min\Big(r_{i}\cdot A_{i},\ \text{clip}(r_{i},1-\epsilon,1+\epsilon)\cdot A_{i}\Big)$ (2)
$-\ \beta\,\mathrm{D}_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Big]$
where $r_{i}=\tfrac{\pi_{\theta}(o_{i})}{\pi_{\theta_{\text{old}}}(o_{i})}$, $\beta$ is a regularization coefficient, preventing excessive deviation from the reference policy during optimization.
@@ 3.2 Spatio-Temporal Rewards of Video MLLM in GRPO
For tasks like video captioning with open-ended outputs, it is impossible to simply compare and determine the gap between the generated caption and the ground truth caption. Therefore, we use a LLM as a “judge" to provide a reward score. In order to reduce the uncertainty of the evaluation criteria for the LLM, we first make the LLM decompose the ground truth and predicted captions into events list. Specifically, we utilize Qwen2.5-72B [38] to extract the events in the description and judge whether the events in a ground truth description can be entailed by the description predicted by the model. We calculate the event recall score as the ratio of events in a ground truth description that are entailed by the predicted description, and set different rewards according to the event recall score:
@@ 3.3 Enhance Spatio-Temporal Perception of Video MLLM through GRPO
We collect diverse video corpus from existing public video datasets for reinforcement learning training to enhance the model’s spatiotemporal perception ability. For the temporal grounding task, we use the training set of Charade - STA [8] (5,338 samples) for training. For the object tracking task, training is conducted on the GoT - 10k [11] dataset, which has 9,335 samples. For the QA and grounding QA tasks, the validation set of NExTGQA  [35] (3,358 samples) is used for training. For video captioning, FIBER-1k [36] (1,000 samples) is adopted for training. For video quality assessment, we use the quality assessment task from VidTAB [17] under the 100-shot setting, with 200 samples for training. Finally, for the training of VideoChat-R1, we perform joint training on three spatio-temporal perception-related tasks: temporal grounding, object tracking, and grounding QA. In total, 18,031 samples are used for training.
@@
Algorithm 1 VideoChat $\mathcal{M}$ with "Clue-Perception"
@@
After enhancing the spatiotemporal perception capability of MLLMs, we can construct a more reliable video chat system. Specifically, after the model answers a user’s question, it can provide relatively accurate clues that support the answer to that question. We can further leverage these clues to improve the accuracy of the responses. Herein, we propose a simple "Clue-Perception" operation: after the model gives the first answer, we re-input the video segments corresponding to the obtained clues into the model at a higher resolution and frame rate, prompting it to answer again. By perceiving more details, the model can generate more accurate responses. Additionally, this operation is also well-suited for long video understanding under conditions of limited computing resources.
@@ 4 Experiments
The main experiments are all conducted based on Qwen2.5-VL-7B [1] (except for the video captioning, for which Qwen2-VL-7B [30] is used).
@@
We employ MVBench [16], Perception Test [22], VideoMME [7] for evaluation of general video understanding. Given that the majority of videos in our training set are short-length, we only use the short subset of VideoMME in testing. For the temporal grounding task, we use the test set of Charade-STA [8] for in-domain testing and the test set of ActivityNet-Grounding [14] as out-domain test data. For the object tracking task, testing is done using the GoT-10k [11] dataset. For the QA and grounding QA tasks, the test set of NExTGQA  [35] is used for testing. And we use Dream-1k [29] and VidTAB-QA [17] for the video captioning and video quality access.
@@ 4.1 Evaluation of VideoChat-R1
Table 1: Results of VideoChat-R1 on various Video Benchmarks. * indicates that the model has suffered from overfitting and is unable to answer the question properly. Since the number of input pixels is fixed during our evaluation, the baseline results are slightly lower than those reported in their origin paper [1].
@@ 4.2 Ablation Studies and Discussions
Method Epochs Training Prompt Test Prompt Charades-STA (in domain) ActivityNet (out domain) VideoMME Think Answer Think Answer mIoU R@0.3 R@0.5 R@0.7 mIoU R@0.3 R@0.5 R@0.7 Short-Avg Vision Experts FlashVTG [2] - - - - - - - 70.3 49.9 - - - - - InternVideo2-6B [32] - - - - - - - 70.0 49.0 - - - - - SG-DETR [9] - - 71.1 52.8 - - - - - MLLMs Qwen2.5-VL-7B - - - 29.0 44.7 24.2 11.1 21.1 28.3 15.8 7.4 71.3 (baseline) - - - 28.1 41.8 23.4 11.1 17.7 22.7 13.4 7.7 71.3 + SFT 1 46.3 63.9 45.0 25.3 20.6 30.2 16.7 7.9 N/A*(-71.3) 3 34.6(+6.5) 51.7 36.3 20.6 17.3(-3.8) 26.1 10.0 3.9 N/A*(-71.3) + GRPO 1 58.7 80.9 67.7 45.4 31.9 46.3 28.8 14.1 72.6 1 59.3(+31.2) 81.7 70.4 46.0 30.7(+13.0) 45.0 27.5 12.9 73.6(+2.3) 3 61.3(+33.2) 83.1 72.8 51.5 34.3(+16.6) 50.4 32.2 16.2 70.9(-0.4)
@@
Table 8: Ablation results of "Clue-Perception". It should be noted that due to our adoption of a lower number of input pixels, the absolute performance is not entirely consistent with that reported for Qwen2.5-VL.
@@
As shown in Table. 8, we compared the performance changes of the model with and without perception enhancement when applying the "Clue-Perception" strategy on two representative long video benchmarks [7, 34]. It is noteworthy that without the use of "Clue-Perception", VideoChat-R1 showed no significant performance improvement over Qwen2.5-VL-7B in long video tasks, which can be attributed to the fact that our training dataset consists entirely of short videos under 1 minute. However, after the application of the "Clue-Perception" operation, VideoChat-R1 demonstrated a significant performance enhancement, indicating that the clues it provides are more accurate and thus revealing the potential of clue-perception in long video understanding. In contrast, due to its insufficient spatiotemporal perception capability, Qwen2.5-VL-7B even exhibited a performance decline after the implementation of the "Clue-Perception" operation.

### v5

@@ Abstract
Reinforcement Learning (RL) benefits Large Language Models (LLMs) for complex reasoning. Inspired by this, we explore integrating spatio-temporal specific rewards into Multimodal Large Language Models (MLLMs) to address the unique challenges of video understanding, such as long-range temporal associations. This paper investigates how rule-based rewards, particularly temporal ones, can improve video reasoning and their generalizability. Our study proposes Reinforcement Fine-Tuning (RFT) as a data-efficient method to enhance video reasoning on specific tasks without sacrificing original capabilities. Through joint RFT on multiple spatio-temporal perception tasks, we developed VideoChat-R1, a powerful Video MLLM. VideoChat-R1 achieves state-of-the-art spatio-temporal perception, demonstrating significant improvements in tasks like temporal grounding (+31.8) and object tracking (+31.2), while also improving general QA benchmarks. The enhanced perception and preserved chat abilities contribute to a more reliable video dialogue system, leading to our “Temporal Clue-driven Reasoning" inference schema. This work provides a foundation for developing robust, real-world video comprehension agents.
@@ 1 Introduction
The integration of reinforcement learning (RL) has notably propelled the capabilities of large language models (LLMs) forward, particularly in complex reasoning. Studies, such as o-series from OpenAI [jaech2024openai] and R1 of DeepSeek [guo2025deepseek], demonstrated that test-time scaling improves model reasoning and such reasoning is elicit-able via only rule-based rewarding. Inspired by these advances, community is interested in tuning multimodal large language models (MLLMs) with vision-related verifications in Group Relative Policy Optimization (GRPO) [guo2025deepseek] manner  [zhou2025r1, segzero, visonr1, deng2025boosting, peng2025lmm, visualrft, r1ov, r1vl, zhan2025visionr1, deng2025openvlthinker]. They primarily consider visual mathematical reasoning and spatial localization.
Regarding video understanding, reasoning is also crucial as some of its core abilties (e.g. its long-range temporal association, fine-grained spatiotemporal understanding based on user queries) are barely addressed via perception or simple analysis. Initial works [r1omni, wang2025timezero, videor1] have validated the superiority of the GRPO algorithm over supervised fine-tuning in some specific video tasks, such as temporal grounding and video question answer. Considering advancing the agent development with real-world video comprehension, we need to investigate and integrate spatio-temporal specific rewards or verification mechanisms directly into the learning of current MLLMs in a scalable manner. This could foster a more tightly coupled, near-closed-loop learning environment, enabling MLLMs to master intricate spatio-temporal skills. While our current work may not fully achieve this ambitious goal, it endeavors to lay foundational groundwork for the community by exploring how to embed such mechanisms effectively. Specifically, we in this paper chart the landscape of how rule-based rewards (especially temporal related ones) work in video understanding, along with systematic evaluations of the its generalizability across diverse video-based reasoning scenarios.
@@
Our study begins with spatio-temporal perception, extending to other popular video tasks (e.g. video caption), forming a practical recipe to enhance video reasoning via proper Reinforcement Fine-Tuning (RFT). We evaluate the robustness of various video verifications through different training and testing setting, also benchmarking their respective data requirements. Moreover, how these rewards interact with each other are also systematically explored. Based on these experiences, we give an intuitive and effective inference schema named “Temporal Clue-driven Reasoning”. Generally, our main findings are as follows.
@@
  • Through joint reinforcement fine-tuning on multiple spatio-temporal perception tasks, we construct VideoChat-R1, a powerful Video MLLM that boasts state-of-the-art spatiotemporal perception capabilities while also taking into account chat abilities. We have also discovered that training on spatio-temporal perception tasks has slightly strengthened the model’s spatio-temporal reasoning abilities. Compared with Qwen2.5-VL-7B, VideoChat-R1 achieves several times the performance improvement in spatiotemporal perception tasks such as temporal grounding (+31.8) and object track (+31.2). At the same time, it also achieves non-trivial improvements on general QA benchmarks, such as VideoMME (+0.9), MVBench (+1.0), and Perception Test (+0.9)
  • The improvement of spatio-temporal perception ability and the preservation of the original chat capability can contribute to a more reliable and efficient video dialogue system. Our VideoChat-R1 can provide reference video segments when answering users’ questions. Meanwhile, we propose to utilize these video segments for “Temporal Clue-driven Reasoning" to further obtain more accurate answers. Our experimental results reveal the potential of the approach that enhances the model’s spatiotemporal perception ability through reinforcement learning for future research in the directions of reliable video dialogue systems and long video understanding.
@@ 2 Related work
Reasoning in MLLMs.
The field LLMs has witnessed significant breakthroughs in reasoning, exemplified by recent works [jaech2024openai, guo2025deepseek]. These developments [shao2024deepseekmath, guo2025deepseek, team2025kimi15] have markedly improved LLMs’ proficiency in tackling complex, multi-step tasks, such as challenging mathematical problems and coding exercises. Extending these RL techniques to multimodal large language models (MLLMs) has seen numerous efforts [zhou2025r1, segzero, visonr1, deng2025boosting, peng2025lmm, visualrft, r1ov, r1vl, zhan2025visionr1, deng2025openvlthinker] focused on leveraging verifiable reward mechanisms to boost visual reasoning performance. However, the application of RL-based strategies to spatio-temporal reasoning within the video domain remains comparatively less explored, with a few studies [wang2025timezero, r1omni, videor1] investigating this area. [wang2025timezero] and  [r1omni] show GRPO’s potential in temporal grounding and sentiment analysis, while,  [videor1] demonstrates how GRPO can facilitate temporal reasoning.
Spatio-Temporal Understanding with MLLMs.
Video understanding heavily relies on spatio-temporal perception. Progress in Video Multimodal Large Language Models (video MLLMs) [li2023videochat, maaz2023video, li2024mvbench, wang2024internvideo2, zhang2024video, wang2025internvideo25, li2024videochatf, videochatonline, qwen25vl] addresses video question answering and captioning in the unfied dialogue format, leveraging LLMs to organize predictions over visual perceptions.  [yu2024merlin] and  [zeng2024timesuite] leverage data augmentation for temporal modeling, yet compromising their general performance.  [yan2024task] improves fine-grained spatio-temporal perception by task preference optimization with solid supervised fine-tuning using a lot of data.
@@ 3 Methodology
We present how to exploit the spatio-temporal rewards to improve MLLMs in video domains using GRPO. Before that, we revisit the Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] first.
3.1 Preliminary
Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] compares groups of candidates responses directly, independent on critic models. Regarding this, it significantly lowers training resources. Formally, with an input query $q$, GRPO initiates by generating a set of $G$ distinct candidate responses, denoted as $o=\{o_{1},\dots,o_{G}\}$, via policy sampling. Each of these candidates is then assigned a score from a predefined reward function, yielding $R_{1},\dots,R_{G}$}. To normalize these scores and ascertain the relative quality of each response, GRPO computes their mean and standard deviation:
$A_{i}=\frac{R_{i}-\mathrm{mean}(\{R_{i}\}_{i=1}^{G})}{\mathrm{std}(\{R_{i}\}_{i=1}^{G})}\text{.}$ (1)
Here $A_{i}$ quantifies the relative quality of the $i$-th candidate response within its group. The objective of GRPO is to encourage the model to prioritize and generate responses that achieve higher relative scores within such groups. Furthermore, the overall training objective incorporates a KL-divergence term $\mathrm{D}_{\mathrm{KL}}(\cdot|\cdot)$, which serves to constrain the optimized policy $\pi_{\theta}$ from diverging excessively from the original MLLM parameters $\pi_{\mathrm{ref}}$.
@@ 3.2 Spatio-Temporal Rewards of Video MLLM in GRPO
For tasks like video captioning with open-ended outputs, it is impossible to simply compare and determine the gap between the generated caption and the ground truth caption. Therefore, we use a LLM as a “judge" to provide a reward score. In order to reduce the uncertainty of the evaluation criteria for the LLM, we first make the LLM decompose the ground truth and predicted captions into events list. Specifically, we utilize Qwen2.5-72B [qwen25] to extract the events in the description and judge whether the events in a ground truth description can be entailed by the description predicted by the model. We calculate the event recall score as the ratio of events in a ground truth description that are entailed by the predicted description, and set different rewards according to the event recall score:
@@ 3.3 Enhance Spatio-Temporal Perception of Video MLLM through GRPO
We collect diverse video corpus from existing public video datasets for reinforcement learning training to enhance the model’s spatiotemporal perception ability. For the temporal grounding task, we use the training set of Charade - STA [charadesta] (5,338 samples) for training. For the object tracking task, training is conducted on the GoT - 10k [got] dataset, which has 9,335 samples. For the QA and grounding QA tasks, the validation set of NExTGQA  [nextgqa] (3,358 samples) is used for training. For video captioning, FIBER-1k [fiber1k] (1,000 samples) is adopted for training. For video quality assessment, we use the quality assessment task from VidTAB [videoeval] under the 100-shot setting, with 200 samples for training. Finally, for the training of VideoChat-R1, we perform joint training on three spatio-temporal perception-related tasks: temporal grounding, object tracking, and grounding QA. In total, 18,031 samples are used for training.
@@
Algorithm 1 VideoChat $\mathcal{M}$ with “Clue-Perception"
@@
After enhancing the spatiotemporal perception capability of MLLMs, we can construct a more reliable video chat system. Specifically, after the model answers a user’s question, it can provide relatively accurate clues that support the answer to that question. We can further leverage these clues to improve the accuracy of the responses. Herein, we propose a simple “Temporal Clue-driven Reasoning" schema: after the model gives the first answer, we re-input the video segments corresponding to the obtained clues into the model at a higher resolution and frame rate, prompting it to answer again. By perceiving more details, the model can generate more accurate responses. Additionally, this operation is also well-suited for long video understanding under conditions of limited computing resources.
@@ 4 Experiments
The main experiments are all conducted based on Qwen2.5-VL-7B [qwen25vl] (except for the video captioning, for which Qwen2-VL-7B [qwen2vl] is used).
@@
We employ MVBench [li2024mvbench], Perception Test [patraucean2023perception], VideoMME [videomme] for evaluation of general video understanding. Given that the majority of videos in our training set are short-length, we only use the short subset of VideoMME in testing. For the temporal grounding task, we use the test set of Charade-STA [charadesta] for in-domain testing and the test set of ActivityNet-Grounding [anetgrounding] as out-domain test data. For the object tracking task, testing is done using the GoT-10k [got] dataset. For the QA and grounding QA tasks, the test set of NExTGQA  [nextgqa] is used for testing. And we use Dream-1k [dream1k] and VidTAB-QA [videoeval] for the video captioning and video quality access.
@@ 4.1 Evaluation of VideoChat-R1
Table 1: Results of VideoChat-R1 on various Video Benchmarks. * indicates that the model has suffered from overfitting and is unable to answer the question properly. Since the number of input pixels is fixed during our evaluation, the baseline results are slightly lower than those reported in their origin paper [qwen25vl].
@@ 4.2 Ablation Studies and Discussions
Method Epochs Training Prompt Test Prompt Charades-STA (in domain) ActivityNet (out domain) VideoMME Think Answer Think Answer mIoU R@0.3 R@0.5 R@0.7 mIoU R@0.3 R@0.5 R@0.7 Short-Avg Vision Experts FlashVTG [cao2024flashvtg] - - - - - - - 70.3 49.9 - - - - - InternVideo2-6B [wang2024internvideo2] - - - - - - - 70.0 49.0 - - - - - SG-DETR [gordeev2024saliency] - - 71.1 52.8 - - - - - MLLMs Qwen2.5-VL-7B - - - 29.0 44.7 24.2 11.1 21.1 28.3 15.8 7.4 71.3 (baseline) - - - 28.1 41.8 23.4 11.1 17.7 22.7 13.4 7.7 71.3 + SFT 1 46.3 63.9 45.0 25.3 20.6 30.2 16.7 7.9 N/A*(-71.3) 3 34.6(+6.5) 51.7 36.3 20.6 17.3(-3.8) 26.1 10.0 3.9 N/A*(-71.3) + GRPO 1 58.7 80.9 67.7 45.4 31.9 46.3 28.8 14.1 72.6 1 59.3(+31.2) 81.7 70.4 46.0 30.7(+13.0) 45.0 27.5 12.9 73.6(+2.3) 3 61.3(+33.2) 83.1 72.8 51.5 34.3(+16.6) 50.4 32.2 16.2 70.9(-0.4)
@@
Table 8: Ablation results of “Clue-Perception". It should be noted that due to our adoption of a lower number of input pixels, the absolute performance is not entirely consistent with that reported for Qwen2.5-VL.
@@
As shown in Table. 8, we compared the performance changes of the model with and without perception enhancement when applying the "Clue-Perception" strategy on two representative long video benchmarks [videomme, longvideobench]. It is noteworthy that without the use of "Clue-Perception", VideoChat-R1 showed no significant performance improvement over Qwen2.5-VL-7B in long video tasks, which can be attributed to the fact that our training dataset consists entirely of short videos under 1 minute. However, after the application of the "Clue-Perception" operation, VideoChat-R1 demonstrated a significant performance enhancement, indicating that the clues it provides are more accurate and thus revealing the potential of clue-perception in long video understanding. In contrast, due to its insufficient spatiotemporal perception capability, Qwen2.5-VL-7B even exhibited a performance decline after the implementation of the "Clue-Perception" operation.